### v1

@@ 1 Introduction
  • Evidence that the stochastic actor-critic technique with prioritized experience replay outperforms non-distributional techniques and classic algorithms, marking the first extensive sim-to-real evaluation for mapless navigation of terrestrial mobile robots using parallel distributional Deep-RL approaches.
A video showcasing the experiments can be accessed at: https://youtu.be/cOVOijEwLUA. Code, Docker image, and pre-trained models are available at: https://github.com/victorkich/Parallel-Turtle-DRL.
@@ 2 Related Works
Traditional Deep-RL models rewards as a single value, but Bellemare et al. [8] proposed modeling it as a probabilistic distribution. This idea was extended to the Soft Actor-Critic (SAC) algorithm by Duan et al. [9], which inspired our approaches.
@@ 3.1 Deep Reinforcement Learning
The objective of standard reinforcement learning is to maximize the expectation of the sum of discounted rewards. The state-action value function is a mathematical model that describes the expected return when taking an action $a$ from a state $s$, and then acting according to the policy $\pi$. This function, known as
@@ 3.2 Parallel Distributional Deterministic RL
The DDPG [14] architecture, proposed as an extension to the DQN, has been a cornerstone in Deep-RL applications of mobile robots in continuous observation spaces [15]. It adopts an actor-critic approach that employs approximation functions to learn continuous space policies.
The main difference between DDPG and D4PG is the use of the distributional Bellman operator to perform the optimization process. This function can be defined as
@@
where $Q_{\pi}(s,a)=\mathbb{E}\ Z_{\pi}(s,a)$ and $Z_{\pi}$ returns a distributional variable. Another important feature of D4PG is the use of the $N$-step return [11] when estimating the TD error, turning the loss function into
@@
Therefore, we have developed a deterministic approach called PDDRL, which is based on the D4PG. Our approach includes an extension of n-step returns and a critic value function modeled as a categorical distribution [8] with an exploration policy added to the actor-network policy. The deterministic policy is denoted by $\mu$ while a noisy action is represented by $\mu^{\prime}$, and the noise process is described as:
@@
where $\mathcal{N}$ is the Ornstein-Uhlenbeck noise method [16].
Because of the non-consistent experimental results obtained by the original D4PG paper [7], we decided to fork our approach and train the agent using a classical and a prioritized replay memory [17]. So, we have an approach called PDDRL which uses the classical replay memory, and another called PDDRL-P that uses prioritized replay memory.
Finally, a target network needs to be established to control the learning and stability. This network is a clone of the actor and critic network with ”soft” updates. The new $\theta^{\prime}$ weights are slowly modified, taking into account the $\tau$ factor:
@@ 3.3 Parallel Distributional Stochastic RL
The SAC architecture [18] was proposed as a stochastic version of the actor-critic deterministic techniques such as the DDPG. The SAC algorithm adopts a bias-stochastic actor-critic structure, which employs approximation functions to learn policies in continuous action spaces. Differently from (2), SAC uses a Bellman function with addition of entropy, which is defined as
$\footnotesize(T_{\pi}Q)(s,a)=r(s,a)+\gamma\mathbb{E}\left[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\ |\ s,a\right].$ (7)
In addition to maximizing the reward, the SAC algorithm aims to increase the entropy of the policy. Entropy is a measure of the unpredictability of a variable, and higher entropy encourages exploration of new states. When a variable remains constant, its entropy is zero, and there is no incentive to modify it. To promote exploration, the algorithm assigns equal probability to actions with the same Q-value and ensures that the policy does not fail due to a single uncertain action in the Q-function approximation.
Similarly to the deterministic case, the main difference between SAC and DSAC is the Bellman operator of DSAC that uses the maximum entropy of (7) and the distributional of (3), creating a distributional soft Bellman defined as
$\footnotesize(T_{\pi}Z)(s,a)=r(s,a)+\gamma\left[Z(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\ |\ s,a\right],$ (8)
where
$\small Z_{\pi}(s,a)=\sum_{t=0}^{\infty}\left[r(s,a)+\gamma-\alpha\log\pi(a_{t+1}|s_{t+1})\ |\ s,a\right].$ (9)
We have proposed a novel method, which we refer to as PDSRL, based on the DSAC algorithm [19]. For a fair comparison, we also implemented an approach (PDSRL) that uses the classical replay memory and another that uses prioritized replay memory (PDSRL-P). A $100000$-step sized replay memory was used for all Deep-RL approaches as well.
@@ 3.4 Network Structure
Figure 4: Moving average of the agent’s reward at each training step for all parallel approaches in each scenario.
@@
3.5 Simulation and Real Scenarios
The simulations of the robot and the scenario use the Gazebo simulator [20]. The connection between the robot and the agents was made using Robot Operating System (ROS). For distributing the agents, multiple instances of the Gazebo simulator are created in parallel and the data acquired are published to the replay buffer. The robot chosen here is the TurtleBot3 Burger version. The real and simulated Turtlebot3 Burger used can be seen in Fig. 3.
Four simulation scenarios were employed in this study. The first scenario represents a navigable area for the robot to move, with the walls being the only obstacle that could cause the robot to collide. In the event of a collision with the wall or any obstacle, a negative reward is issued for the action, and the current episode is terminated. The second scenario features four fixed cylinder-shaped obstacles with a radius of 25 centimeters each. The third and fourth scenarios were built in order to create more challenging paths for the robot to reach the final goal. The third scenario includes a “U”-shaped object that presents the agent with two possible trajectories of equal cost, with the obstacle creating a dead-end in the middle. The fourth scenario is asymmetrical and more complex, requiring the intelligent agent to develop better strategies to avoid collisions. Both real and simulated scenarios are appropriately labeled in Fig. 3.
@@ 3.6 Reward Function
4 Experimental Results
We evaluated the navigation ability and spatial generalization of all approaches in both simulation and real-world scenarios. After the training phase, consisting of $30000$ steps for the first scenario and $200000$ steps for the others, we analyzed the moving average of rewards (Fig. 4).
The results showed that stochastic distributional approaches outperformed others in all scenarios, despite a slower start. DDPG had similar reward means to the top approaches. In more complex scenarios, approaches without prioritized memory had slower knowledge acquisition.
Each approach was evaluated over $100$ episodes with predetermined target coordinates, as shown in Fig. 4.
@@ 4 Experimental Results
Fig. 5 shows PDSRL and PDSRL-P had smoother trajectories. DDPG showed instability with repeated collisions in Scenario 3. Prioritization reduced collisions in Scenario 4, but instability remained.
Figure 7: The behavior of each parallel approach was evaluated by conducting 12 navigation trials in each real scenario. The lines illustrate the paths taken by the agents, with each agent given 3 attempts to capture each target.
Real environment experiments used target points similar to simulations. Fig. 7 shows trajectories in all real scenarios. PDSRL-P and DDPG-P achieved $100\%$ accuracy in real environments. Performance differences between PDDRL and PDDRL-P were significant in scenarios 3 and 4. Real scenarios introduced delays, affecting performance.
Figure 8: Behavior of the best classified parallel approaches evaluated in the extra real scenario over 12 navigation trials. The lines illustrate the paths taken by the agents, with each agent given 3 attempts to capture each target.

### v2

@@ 1 Introduction
  • Evidence that the stochastic actor-critic technique with prioritized experience replay outperforms non-distributional techniques and classic algorithms, marking the first extensive sim-to-real evaluation for mapless navigation of terrestrial mobile robots using parallel distributional Deep-RL approaches.  [^1] https://youtu.be/cOVOijEwLUA [^2] https://github.com/victorkich/Parallel-Turtle-DRL
@@ 2 Related Works
Traditional Deep-RL models rewards as a single value, but Bellemare et al. [8] proposed modeling it as a probabilistic distribution. This idea was extended to the new implementation of Soft Actor-Critic (SAC) algorithm by Duan et al. [9], which inspired our approaches.
@@ 3.1 Deep Reinforcement Learning
The objective of standard reinforcement learning is to maximize the expectation of the sum of discounted rewards. The state-action value function is a mathematical model that describes the expected return when taking an action $a$ from a state $s$, and then acting according to the policy $\pi$ [14]. This function, known as
@@ 3.2 Parallel Distributional Deterministic RL
The DDPG architecture, as proposed by [15] and extended in D4PG [7], serves as a cornerstone in Deep-RL applications for mobile robots in continuous observation spaces [16]. It adopts an actor-critic framework that utilizes approximation functions to learn policies in continuous spaces.
The primary distinction between DDPG and D4PG lies in the latter’s incorporation of the distributional Bellman operator, which is crucial for the optimization process. This operator is defined as follows:
@@
where $Q_{\pi}(s,a)=\mathbb{E}\ Z_{\pi}(s,a)$ and $Z_{\pi}$ returns a distributional variable, specifically a categorical distribution.
In the D4PG framework, the categorical distribution models the output of the critic network, which predicts a vector of probabilities over predefined reward bins, each corresponding to a range of potential reward values. This distribution allows the network to estimate a full probability distribution of expected returns, rather than a single expected value, capturing the variability in possible outcomes:
$\text{Categorical}(Z_{\pi}(s,a))=\left[\ p_{1},p_{2},\dots,p_{k}\ \right],$ (4)
where $p_{i}$ is the probability of the return falling into the $i$-th bin and $k$ is the number of bins. The training of the critic involves minimizing the divergence between the predicted and target distributions, enhancing the policy robustness by better representing the uncertainties in dynamic environments.
Furthermore, D4PG introduces the concept of $N$-step returns to estimate the Temporal Difference (TD) error, expressed as:
@@
In our research, we have developed an approach called PDDRL based on D4PG, incorporating an extension of $N$-step returns and a critic value function modeled as a categorical distribution [8]. The deterministic policy is denoted by $\mu$, while noisy actions are represented by $\mu^{\prime}$, where the noise process involves:
@@
and $\mathcal{N}$ follows the Ornstein-Uhlenbeck process [17].
Given the inconsistent experimental results reported in the original D4PG study [7], our research has led us to develop two variants: PDDRL, which utilizes classical replay memory, and PDDRL-P, employing prioritized replay memory [18].
Finally, a target network needs to be established to enhance learning stability. This network is a duplicate of the actor and critic networks but uses “soft” updates. The weights $\theta^{\prime}$ of the target network are gradually adjusted according to the factor $\tau$, as described by the equation:
@@ 3.3 Parallel Distributional Stochastic RL
The SAC architecture [19], proposed as a stochastic counterpart to deterministic actor-critic methods like DDPG, employs approximation functions to learn policies in continuous action spaces. SAC introduces a Bellman operator enhanced by entropy addition, aiding in exploration and policy optimization:
$(T_{\pi}Q)(s,a)=r(s,a)+\gamma\mathbb{E}\left[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\ |\ s,a\right].$ (8)
To foster robust exploration, SAC emphasizes maximizing both reward and the entropy of the policy, thereby promoting action diversity and discouraging premature policy convergence. The algorithm assigns equal probability to actions yielding similar Q-values and mitigates failure risks in the Q-function approximation due to uncertain actions.
Building on this, DSAC [20] merges the maximum entropy framework of SAC with the distributional approach of DDPG, leading to a hybrid Bellman operator:
$(T_{\pi}Z)(s,a)=r(s,a)+\gamma\left[Z(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\ |\ s,a\right],$ (9)
which integrates both stochastic and distributional components, further described as:
$Z_{\pi}(s,a)=\sum_{t=0}^{\infty}\left[r(s,a)+\gamma-\alpha\log\pi(a_{t+1}|s_{t+1})\ |\ s,a\right].$ (10)
In our research, the newly proposed PDSRL methodology builds upon DSAC, incorporating the strategic use of N-step returns and soft updates—features previously detailed in Section 3.2 under PDDRL—to enhance both prediction accuracy and stability.
We implemented two variants, PDSRL using classical replay memory, and PDSRL-P with prioritized replay memory, both featuring a $100000$-step sized replay memory for consistency across all Deep-RL approaches.
@@ 3.4 Network Structure
Figure 4: Moving average of the agent’s reward at each training step for all parallel approaches in each scenario. Scenarios organized by following the respectively order, from left to right and from top to bottom: 1, 2, 3, and 4.
@@ 4 Experiments
In this section, we present the experimental evaluation of our proposed approaches. The experiments are designed to assess the performance of the methods in both simulated and real-world environments. We describe the setup of the simulated and real scenarios used for training and testing, followed by a detailed analysis of the experimental results.
4.1 Simulation and Real Scenarios
The simulations of the robot and the scenario use the Gazebo Simulator. The connection between the robot and the agents was made using Robot Operating System (ROS). For distributing the agents, multiple instances of the Gazebo simulator are created in parallel and the data acquired are published to the replay buffer. The robot chosen here is the TurtleBot3 Burger version. The real and simulated Turtlebot3 Burger used can be seen in Fig. 3.
Four simulation scenarios were employed in this study. The first scenario represents a navigable area for the robot to move, with the walls being the only obstacle that could cause the robot to collide. In the event of a collision with the wall or any obstacle, a negative reward is issued for the action, and the current episode is terminated. The second scenario features four fixed cylinder-shaped obstacles with a radius of 25 centimeters each. The third and fourth scenarios were built in order to create more challenging paths for the robot to reach the final goal. The third scenario includes a “U”-shaped object that presents the agent with two possible trajectories of equal cost, with the obstacle creating a dead-end in the middle. The fourth scenario is asymmetrical and more complex, requiring the intelligent agent to develop better strategies to avoid collisions. Both real and simulated scenarios are labeled in Fig.3.
4.2 Experimental Results
We evaluated the navigation ability and spatial generalization of all approaches in both simulation and real-world scenarios. After the training phase, consisting of $30000$ steps for the first scenario and $200000$ steps for the others, we analyzed the moving average of rewards (Fig. 4). The results showed that stochastic distributional approaches outperformed others in all scenarios, despite a slower start. DDPG had similar reward means to the top approaches. Each approach was evaluated over $100$ episodes with predetermined target coordinates.
@@ 4.2 Experimental Results
Fig. 5 shows PDSRL and PDSRL-P had smoother trajectories. DDPG showed instability with repeated collisions in Scenario 3. Prioritization reduced collisions in Scenario 4, but instability remained. Real environment experiments used target points similar to simulations. Fig. 7 shows trajectories in all real scenarios. PDSRL-P and DDPG-P achieved $100\%$ accuracy in real environments. Performance differences between PDDRL and PDDRL-P were significant in scenarios 3 and 4. Real scenarios introduced delays, affecting performance.
Figure 7: (Top) The behavior of each parallel approach was evaluated by conducting 12 navigation trials in each real scenario. The lines illustrate the paths taken by the agents, with each agent given 3 attempts to capture each target. (Bottom) The behavior of the best classified parallel approaches evaluated in the extra real scenario. Organized by the following order: DDPG, DDPG-P, PDDRL, PDSRL, and PDSRL-P.